OpenStudy (anonymous):

write radical expression in exponential form: 8 divided by 7th root of x to 15th power

jhonyy9 (jhonyy9):

so like an example if squarrot of x is equal x on exponent 1/2 where the denominator 2 is the indice of root hope so much that now using this example you will can solveing your exercise easy ok ? good luck bye

OpenStudy (anonymous):

is it 8x7/15?

jhonyy9 (jhonyy9):

so if squarroot of x is x on exponent 1/2 than 7th root of x on 15th power how will be right ?

OpenStudy (anonymous):

I don't understand what you are saying

jhonyy9 (jhonyy9):

how you write 7th root of x ?

OpenStudy (anonymous):

sq rt. sign with x underneath and 7 outside to left

jhonyy9 (jhonyy9):

\[\sqrt[7]{x}\]

OpenStudy (anonymous):

yes but x has exponent of 15

jhonyy9 (jhonyy9):

so than how will rewrite you this using exponents ?

OpenStudy (anonymous):

x15/7

jhonyy9 (jhonyy9):

ok i understand it but i wann you understand it step by step ok ?

OpenStudy (anonymous):

what am I missing?

jhonyy9 (jhonyy9):

so when there are 7th root of x^15 how can you simplifie it ? or what can you factoriz out from root ?

OpenStudy (anonymous):

oh- what can be factored out of 7-nothing?

jhonyy9 (jhonyy9):

so check it please how can you rewrite it x on exponent 15 using exponent of 7 ?

OpenStudy (anonymous):

no idea

jhonyy9 (jhonyy9):

why ? this is very very easy

OpenStudy (anonymous):

for u it is. not sure what u are asking

jhonyy9 (jhonyy9):

when 15=7+7+1

jhonyy9 (jhonyy9):

or for your easy understanding x^15 =x^7 *x^7 *x right ?

OpenStudy (anonymous):

but I don't understand what u want me do with the factoring. my answer should have 7/15 or 15/7 in it and if you factor out it won't

OpenStudy (anonymous):

yes-see I need the easy explanation

jhonyy9 (jhonyy9):

so x^15 = (x^2)^7 *x ok ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than how you calcule 7th root of x^15 ?

OpenStudy (anonymous):

why do I need to calculate 7th rt of x^15? Don't I have to eliminate the denominator and multiply the numerator by that to get answer?

jhonyy9 (jhonyy9):

ok so you need to write the 8/(7th root of x^15) in a simplified form ?

OpenStudy (anonymous):

my question states in exponential form but I guess it is simplified form. My choices have no denom. only numerator 8x with ^either -15/7 or -7/15 or 7/15 or 15/7. I think it is 8x^7/15

jhonyy9 (jhonyy9):

so the 7th root of x on power 15 how you can write it using exponents ?

OpenStudy (anonymous):

15/7?

jhonyy9 (jhonyy9):

so than will be x^(15/7) yes ?

OpenStudy (anonymous):

yes but I have to multiply x^15/7 by x^7/15 to remove from denom.

jhonyy9 (jhonyy9):

so than your exercise will be 8 -------- x^(15/7) ok ?

OpenStudy (anonymous):

but I can't leave that in the denomin

jhonyy9 (jhonyy9):

so when there are x^(-1) = ?

OpenStudy (anonymous):

1/x

jhonyy9 (jhonyy9):

yes ,right

OpenStudy (anonymous):

what is the point you r making

jhonyy9 (jhonyy9):

so than from this resulted that your exercise will be ???

OpenStudy (anonymous):

I thought it would be 8x^7/15

jhonyy9 (jhonyy9):

so there are 8 -------- = ? x^(15/7) using this example of 1/x = x^(-1)

OpenStudy (anonymous):

do I not multiply denom by x ^7/15 to eliminate it and then multip. numerator by same number?

jhonyy9 (jhonyy9):

so this you need to do it when you need to rationalize the denominator ,but here now you need to rewrite this exercise in a simplified form using exponents ok ?

OpenStudy (anonymous):

will if the answer doesn't have a denominator then I am also rationalizing aren't I?

jhonyy9 (jhonyy9):

yes this is right the answer not will have denominator because how in case of 1/x using exponent with sign minus you can rewrite it like x on exponent minus 1 so the same you need to do in this case of this exercise too ok ?

OpenStudy (anonymous):

\[8x \frac{ -15 }{ 7 }\] is the answer